# Physics of droplet regulation in biological cells

### v2

@@ C Interface properties of binary liquids
We study a flat interface described by the equilibrium profile $\phi_{\mathrm{eq}}(x)$, where the $x$–coordinate is perpendicular to the interface. The profile $\phi_{\mathrm{eq}}(x)$ connects two phases, $\phi_{\mathrm{eq}}(x\rightarrow-\infty)=\phi^{(0)}_{\mathrm{out}}$ and $\phi_{\mathrm{eq}}(x\rightarrow\infty)=\phi^{(0)}_{\mathrm{in}}$, and exhibits a constant chemical potential $\bar{\mu}_{\mathrm{eq}}$,
@@
Using equations (C.4), (C.2), and (C.7), we again find equation (C.6), showing that the mechanical and thermodynamic definitions of $\gamma$ agree.
@@ D Droplets in two dimensions
Droplet growth is still described by equation (2.36), but the diffusive fluxes $\boldsymbol{j}$ are more difficult to determine in two dimensions since steady state concentration profiles diverge logarithmically. As a result, one must introduce a cut-off length in numerical simulations [140] or use asymptotic matching [478, 479]. The latter approach shows that Ostwald ripening is suppressed [478] and that droplets drift in gradients [479], similar to their 3D counterparts. The diffusivity $D_{\mathrm{drop}}$ associated with Brownian motion also exhibits a logarithmic scaling. For a droplet of radius $R$ embedded in a membrane of thickness $h$, the Saffman–Delbrück model predicts [480]
@@ E Derivation of equilibrium stress tensor
Canceling terms including $v_{\alpha}$ due to the divergence theorem and integrating by parts again results in
@@ F Transition state theory
The additional requirement of symmetry between reactants and products, i.e., that swapping reactants and products must turn $s$ into $-s$, implies $\alpha=\frac{1}{2}$, leading to equations (5.4) and (5.5).

### v3

@@ C Interface properties of binary liquids
We study a flat interface described by the equilibrium profile $\phi_{\mathrm{eq}}(x)$, where the xx–coordinate is perpendicular to the interface. The profile $\phi_{\mathrm{eq}}(x)$ connects two phases, $\phi_{\mathrm{eq}}(x\rightarrow-\infty)=\phi^{(0)}_{\mathrm{out}}$ and $\phi_{\mathrm{eq}}(x\rightarrow\infty)=\phi^{(0)}_{\mathrm{in}}$, and exhibits a constant chemical potential $\bar{\mu}_{\mathrm{eq}}$,
@@
Using equations (C.4), (C.2), and (C.7), we again find equation (C.6), showing that the mechanical and thermodynamic definitions of $\gamma$ agree [139].
@@ D Droplets in two dimensions
Droplet growth is still described by equation (2.36), but the diffusive fluxes $\boldsymbol{j}$ are more difficult to determine in two dimensions since steady state concentration profiles diverge logarithmically. As a result, one must introduce a cut-off length in numerical simulations [149] or use asymptotic matching [531, 532]. The latter approach shows that Ostwald ripening is suppressed [531] and that droplets drift in gradients [532], similar to their 3D counterparts. The diffusivity $D_{\mathrm{drop}}$ associated with Brownian motion also exhibits a logarithmic scaling due to a hydrodynamic coupling mediated via the surrounding liquid. For a droplet of radius $R$ embedded in a membrane of thickness $h$, the Saffman–Delbrück model predicts [533]
@@ E Derivation of equilibrium stress tensor
The divergence theorem applied to $\rho v_{\alpha}^{2}$ implies terms involving $v_{\alpha}$ cancel. Integrating by parts then yields
@@ F Transition state theory
The additional requirement of symmetry between reactants and products, i.e., that swapping reactants and products must turn $s$ into s-s, implies $\alpha=\frac{1}{2}$, leading to equations (5.4) and (5.5).